A. Find the slope, x-intercept and y-intercept form of the following equations a) 5x + 2y =-10 b) 13y -2x = 3 c) 25𝑦 + 31π‘₯ βˆ’ 18 = 10𝑦 d) βˆ’3π‘₯ + 4𝑦 βˆ’ 10 = 7π‘₯ βˆ’ 2𝑦 + 50


1
Expert's answer
2021-08-02T14:39:49-0400

"\text{To get the slope of an equation, we make y the subject of the formula, we set y = 0}\\\text{to get the x-intercept and we set x = 0 to get the y-intercept}\\1. \text{From question a, $y = \frac{-5x}{2}-5$, therefore the gradient(m) =$-\frac{5}{2}$ and the y-intercept}\\\text{is -5. }\\\text{Let y = 0, $x = \frac{-10}{5}=-2$ }
\\2. \text{ From question b, $y = \frac{2x}{13}+\frac{3}{13}$, therefore the gradient(m) =$\frac{2}{13}$ and the y-intercept}\\\text{is $\frac{3}{13}$ }\\\text{Let y = 0, $x = -\frac{3}{2}$ }
\\3. \text{ From question c, $y = -\frac{31x}{15}+\frac{18}{15}$, therefore the gradient(m) =$-\frac{31}{15}$ and the y-intercept}\\\text{is $\frac{18}{15}$ }\\\text{Let y = 0, $x = \frac{18}{31}$ }

\\4. \text{ From question d, $y = \frac{10x}{6}+10$, therefore the gradient(m) = $\frac{10}{6}$ and the y-intercept}\\\text{is 10 }\\\text{Let y = 0, x = -6 }"